OpenStudy (anonymous):

A video game sets the points needed to reach the next level based on the function g(x) = 7(3)x − 1, where x is the current level. The hardest setting promises to multiply the points needed in each level according to the function h(x) = 4x. How many points will a player need on the hardest setting of level 4?

OpenStudy (anonymous):

Is g(x) supposed to be 21x - 1 or \[g(x)=7(3)^x-1\]

OpenStudy (anonymous):

the second one @peachpi

OpenStudy (anonymous):

You need to find g(h(x)).

OpenStudy (anonymous):

More specifically g(h(4))

OpenStudy (anonymous):

g(h(x)) = 7(3)^(4x) - 1 = 7(81)^x - 1